# DINO in the Room: Leveraging 2D Foundation Models for 3D Segmentation

### v1

@@ Appendix B Qualitative Results
Figures 4, 5, 6, 7 and 8 highlight the rich semantic features distilled from DINOv2. Points belonging to the same semantic class exhibit similar colors, reflecting alignment along the first three principal axes. For instance, in Figs. 4 and 5, objects such as tables, chairs, books, and walls consistently map to distinct colors across the scene. These results show that the model captures fine-grained semantic representations of 3D scenes without requiring semantic labels during training. Furthermore, Figs. 4 and 5 reveal that objects like chairs and tables are segmented into finer subparts in feature space, surpassing the granularity of the dataset annotations. This semantically rich feature representation significantly enhances the model’s performance for semantic segmentation fine-tuning, as is evident by the quantitative results in the main paper.
@@
Figures 9, 10 and 11 compare the predictions of DITR and PTv3, with a focus on points where DITR benefits from DINOv2 features.
@@
Figure 9: DITR and PTv3 [58] segmentation for ScanNet [8]. The wheels of the cart are barely visible in the 3D point cloud, hence likely causing PTv3 to misclassify it as a shelf. The bottom-right visualization illustrates the differences between PTv3 and DITR predictions. Points where both methods are correct or both are incorrect are shown in black. Points correctly predicted by only one method are color-coded: green for DITR and red for PTv3.

### v2

@@ Appendix C Qualitative Results
Figs. 4, 5, 6, 7 and 8 highlight the rich semantic features distilled from DINOv2. Points belonging to the same semantic class exhibit similar colors, reflecting alignment along the first three principal axes. For instance, in Figs. 4 and 5, objects such as tables, chairs, books, and walls consistently map to distinct colors across the scene. These results show that the model captures fine-grained semantic representations of 3D scenes without requiring semantic labels during training. Furthermore, Figs. 4 and 5 reveal that objects like chairs and tables are segmented into finer subparts in feature space, surpassing the granularity of the datasets’ annotations, which this model does not use. This semantically rich feature representation significantly enhances the model’s performance when used as initialization before fine-tuning for semantic segmentation, as is evident by the quantitative results in the main paper.
@@
Figs. 9, 10 and 11 compare the predictions of DITR and PTv3, with a focus on points where DITR benefits from DINOv2 features.
@@
Figure 9: DITR and PTv3 [62] segmentation for ScanNet [10]. The wheels of the cart are barely visible in the 3D point cloud, likely causing PTv3 to misclassify it as a shelf. By contrast, DITR can leverage semantically rich image features to correctly identify this cart. The bottom-right visualization illustrates the differences between PTv3 and DITR predictions. Points where both methods are correct or both are incorrect are shown in black. Points correctly predicted by only one method are color-coded: green for DITR and red for PTv3.